A decade-old study found that the proportion of high school seniors who felt that "getting rich" was an important personal goal was

66%

. Suppose that we have reason to believe that this proportion has changed, and we wish to carry out a hypothesis test to see if our belief can be supported. State the null hypothesis

H0

and the alternative hypothesis

H1

that we would use for this test.
Answer #1

SOLUTION:-

Here, we have to determine the null and alternative hypothesis:-

H0 : p=66% =66/100= 0.66

Ha : p not= 66% =66/100 0.66 for two tailed test

Where,

H0 is the null hypothesis.

Ha is the alternative hypothesis.